# Polarized QED Cascades over Pulsar Polar Caps

### v1

@@ 3 Photon polarization
(iii) Super-SR. For $e^{\pm}$ with both large pitch angles and high energies ($\theta_{0}=10^{-3}$ and $\varepsilon_{0}=1$ TeV, $2\theta_{d}\approx 1.7\times 10^{-8}$), they are likely to emit photons on timescales shorter than the gyration period, simultaneously losing most of their energies due to radiation reaction. In this super-SR regime, photon emission concentrates near the initial $e^{\pm}$ position on the $+\bm{e}_{y}$ axis, as shown in the bottom-left panel of Fig. 4. Consequently, the emitted photons exhibit a distinct linear polarization directed perpendicular to the plane of magnetic field lines, i.e., $\overline{\xi}_{3}<0$. This is markedly different from the CR and SR regimes. The polarization degree of super-SR photons decreases to zero at high energy, in agreement with the theoretical expectation for $\chi_{e}>1$. However, as we will see below, the existence of this super-SR regime is unlikely for $B_{0}\gtrsim 4\times 10^{11}$ G. This is because a relatively large pitch angle corresponds to newly created $e^{\pm}$ pairs, which typically do not possess sufficient energy to emit in the super-SR regime. Consequently, this unique radiation regime will not be further detailed in this work.
@@ 4.1 S-type QED cascades
As the QED cascade develops, the proportion of SR photons grows, eventually surpassing CR photons. This trend is evident in the temporal evolution of the number ratio of SR photons to CR photons, $N_{\gamma}^{\rm SR}/N_{\gamma}^{\rm CR}$ (see the middle panel of Fig. 5). The growth rate of $N_{\gamma}^{\rm SR}/N_{\gamma}^{\rm CR}$ exhibits significant variability with the initial electron energy $\varepsilon_{0}$, increasing much more rapidly with a larger $\varepsilon_{0}$. At the final simulation time $t=200~{}\rm ns$, $N_{\gamma}^{\rm SR}\approx N_{\gamma}^{\rm CR}$ for $\varepsilon_{0}=2$ TeV, while $N_{\gamma}^{\rm SR}\gg N_{\gamma}^{\rm CR}$ for both $\varepsilon_{0}=3$ TeV and 4 TeV. The weakly polarized SR photons emitted later contribute to the decrease in the average photon polarization $\overline{\xi}_{3}$, depending also on the initial electron energy $\varepsilon_{0}$ (see the right panel of Fig. 5). For $\varepsilon_{0}=2$ TeV, the effects of pair production are relatively weak, leading to a decrease in average photon polarization from 30% to 15%. For lower energies, $\overline{\xi}_{3}$ rapidly decreases until it converges to a weak SR photon polarization of about 8%.
@@ 4.2 A-type QED cascade
In the top-left panel of Fig. 6, we report the emitted photons in the $\theta$-$\varepsilon_{e}$ space under three curvature radii at the time $t=300$ ns. Unlike the S-type cascade, the low-energy $e^{\pm}$ pairs created here are accelerated to higher energies by the parallel electric field. Consequently, in addition to CR photons from initial electrons ($\theta<10^{-8}$) and SR photons from created $e^{\pm}$ pairs ($\theta>10^{-4}$), we observe additional SR photons emitted by the accelerated $e^{\pm}$ pairs with small pitch angles ($10^{-8}<\theta<10^{-4}$). As in the S-type cascade, super-SR photons are notably absent in all cases.
@@ 4.3 Photon polarization effect on pair production
Figure 7: Left panels: the temporal evolution of the relative difference of the positron number $\Delta_{+}=(N_{+}-\overline{N}_{+})/\overline{N}_{+}$, where $N_{+}$ and $\overline{N}_{+}$ are the number of generated positrons with and without considering the photon polarization, respectively. Right panels: the QED parameter $\chi_{\gamma}$ of decayed photons with and without the photon polarization. Top panel: S-type cascade with $\varepsilon_{0}=3$ TeV and other parameters the same as Fig. 5. Bottom panel: A-type cascade with $\rho=10$ km and other parameters the same as Fig. 6.
@@
In this subsection, we investigate the influence of photon polarization on pair production. According to Eq. (2), the photon polarization can affect the pair production rate by as much as 30%. When applying Eq. (2), one should pay particular attention to the selected basis for the photon Stokes vector $\bm{\xi}$. The two basis vectors $\bm{e}_{1}$ in Eq. (1) and $\bm{e}_{1}^{\prime}$ of pair production in Eq. (2) are mutually perpendicular for CR. This arises from the fact that the acceleration direction of $e^{\pm}$ in CR is mainly parallel to the plane of magnetic field lines, while for newly created $e^{\pm}$ pairs, the acceleration direction is mainly perpendicular to it. Consequently, this results in $\xi_{3}\approx-\xi_{3}^{\prime}$. From Eq. (2), the specific polarization of CR photons leads to a higher pair production rate compared to unpolarized photons. This finding starkly contrasts with results obtained in laser fields, where the photon polarization effect was observed to suppress pair production (Wan et al., 2020; Seipt et al., 2021; Song et al., 2021b). In fact, in laser fields, highly polarized photons are emitted via a SR-like mechanism, leading to $\xi_{3}\approx\xi_{3}^{\prime}$.
@@
The relative difference in the positron number $\Delta_{+}=(N_{+}-\overline{N}_{+})/\overline{N}_{+}$ between polarized and unpolarized results is reported in Fig. 7, where $N_{+}$ and $\overline{N}_{+}$ are the number of generated positrons with and without considering the photon polarization, respectively. Fig. 7 reveals that the number of created positrons is enhanced by more than 10% at the initial stage of the cascade due to the high polarization of CR photons. In the S-type cascade, at later times the relative difference narrows to less than 5%, even though the polarization of CR photons remains high. The comparison of the $\chi_{\gamma}$ of decayed photons with versus without photon polarization (see the right column of Fig. 7) reveals that for the unpolarized case, $\chi_{\gamma}$ at photon decay is slightly larger than in the polarized case. In fact, unpolarized photons with the same energy as polarized photons posses smaller decay rates and travel longer distances, thus acquiring larger pitch angles, and consequently slightly larger $\chi_{\gamma}$ than polarized photons before decaying into pairs. In the A-type cascade, photon polarization has a similar effect on pair production, as illustrated in the bottom panel of Fig. 7. The relative difference $\Delta_{+}$ can exceed 10% at the initial stage and is also reduced to about 5% with the development of the QED cascade. Similar to the S-type cascade, a slight difference in $\chi_{\gamma}$ is observed between polarized and unpolarized results. At a later stage, a key factor to the reduction in the pair yield difference between polarized and unpolarized photons is the domination of unpolarized SR photons over polarized CR photons.
@@
Figure 8: From top to bottom panel: Spatial distributions of the parallel electric field $E_{x}$ and the normalized positron density $n_{+}/|n_{\rm GJ}|$. Emitted $\gamma$ photons in the $\theta$-$\varepsilon_{e}$ space. Temporal evolution of the photon polarization $\overline{\xi}_{3}$ in relation to the photon energy $\varepsilon_{\gamma}$.
@@ 4.4 Polarized QED-PIC simulations
In our final set of simulations, we employ first-principles 1D (1D geometry but fully 3D particle momentum) QED-PIC simulations in the corotating frame using the YUNIC code (Song et al., 2021a). YUNIC is a relativistic PIC code that includes polarized QED modules and has been previously utilized for studying polarized QED cascades in laser-plasma interactions (Song et al., 2021b; Song et al., 2022). Notably, the electric field in these simulations is inhomogeneous and self-consistently solved. This is a crucial distinction from the single-particle simulations, where the electric field is not evolving. Here the electric field is excited by the outflow of charged $e^{\pm}$ plasmas and subsequently screened by the created dense $e^{\pm}$ plasmas, all of which can be captured by the QED-PIC simulation. Our simulation setup is similar to the pioneering work (Timokhin, 2010). There are several improvements in our QED-PIC code, in which a standard Monte-Carlo QED method developed in recent years by the laser-plasma community is adopted (Elkina et al., 2011; Gonoskov et al., 2015; Ridgers et al., 2014; Tamburini et al., 2017; Montefiori & Tamburini, 2023), and enriched with the polarization effects introduced in section 2.
@@
With the propagation of the initial flow of $e^{\pm}$ outside the computational box, a parallel electric field is excited, linearly increasing towards the pulsar surface (left boundary) to accelerate $e^{\pm}$ to high energies (see the top panel of Fig. 8). When the QED cascade starts, the density of created $e^{\pm}$ plasma near the pulsar surface can reach 30 times the GJ density at $t=800$ ns, as shown in the second panel from the top of Fig. 8. As expected, the QED cascade creates large amounts of charges that screen the electric field and close the vacuum gap. The $\theta$-$\varepsilon_{e}$ space of photons at three different instants is displayed in the third panel from top of Fig. 8. Similar to the single-particle simulation results, CR photons have high energies and small pitch angles, while SR photons exhibit opposite features with lower energies and larger pitch angles than CR photons. The transition from primarily having CR photons to SR photons becoming dominant over time is evident. The average photon polarization $\overline{\xi}_{3}$ as a function of the emitted photon energy is shown in the bottom panel of Fig. 8, with the net polarization of low-energy photons gradually reducing due to the growth of the number of weakly polarized SR photons. An additional insight from the PIC simulation is that high-energy photons are also depolarized at a later stage. This is due to the strong induced electric field, which can accelerate a substantial fraction of secondary $e^{\pm}$ to high energies. Before the electric field is screened, these accelerated $e^{\pm}$ emit high-energy photons while remaining in the SR regime. Despite their high energy, these SR photons are comparatively weakly polarized compared to CR photons. With the inclusion of photon polarization effects, we observe an enhancement of approximately 5% in the $e^{\pm}$ yield compared to the unpolarized results.
@@ Acknowledgements
This work was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (Grants No. XDA25050300 and No. XDA25010300), the National Key R&D Program of China (2021YFA1601700).

### v2

@@ 3 Photon polarization
(iii) Super-SR. For $e^{\pm}$ with both large pitch angles and high energies ($\theta_{0}=10^{-3}$ and $\varepsilon_{0}=1$ TeV, $2\theta_{d}\approx 1.7\times 10^{-8}$), they are likely to emit photons on timescales shorter than the gyration period, simultaneously losing most of their energies due to radiation reaction. In this super-SR regime, photon emission concentrates near the initial $e^{\pm}$ position on the $+\bm{e}_{y}$ axis, as shown in the bottom-left panel of Fig. 4. Consequently, super-SR photons exhibit a distinct linear polarization directed along $\bm{e}_{z}$, i.e., the initial transverse acceleration, resulting in $\overline{\xi}_{3}<0$ in the $({\bm{e}}_{y},{\bm{e}}_{z},{\bm{e}}_{x})$ basis. This is markedly different from the CR and SR regimes. The polarization degree of super-SR photons decreases to zero at high energy, in agreement with the theoretical expectation for $\chi_{e}>1$. However, as we will see below, the existence of this super-SR regime is unlikely for $B_{0}\gtrsim 4\times 10^{11}$ G. This is because a relatively large pitch angle corresponds to newly created $e^{\pm}$ pairs, which typically do not possess sufficient energy to emit in the super-SR regime. Consequently, this unique radiation regime will not be further detailed in this work.
@@ 4.1 S-type QED cascades
As the QED cascade develops, the proportion of SR photons grows, eventually surpassing CR photons. This trend is evident in the temporal evolution of the number ratio of SR photons to CR photons, $N_{\gamma}^{\rm SR}/N_{\gamma}^{\rm CR}$ (see the middle panel of Fig. 5). The growth rate of $N_{\gamma}^{\rm SR}/N_{\gamma}^{\rm CR}$ exhibits significant variability with the initial electron energy $\varepsilon_{0}$, increasing much more rapidly with a larger $\varepsilon_{0}$. At the final simulation time $t=200~{}\rm ns$, $N_{\gamma}^{\rm SR}\approx N_{\gamma}^{\rm CR}$ for $\varepsilon_{0}=2$ TeV, while $N_{\gamma}^{\rm SR}\gg N_{\gamma}^{\rm CR}$ for $\varepsilon_{0}=4$ TeV. The weakly polarized SR photons emitted later contribute to the decrease in the average photon polarization $\overline{\xi}_{3}$, depending also on the initial electron energy $\varepsilon_{0}$ (see the right panel of Fig. 5). For $\varepsilon_{0}=2$ TeV, the effects of pair production are relatively weak, leading to a decrease in average photon polarization from 30% to 15%. For larger initial energies, $\overline{\xi}_{3}$ rapidly decreases until it converges to a weak SR photon polarization of about 8%.
@@ 4.2 A-type QED cascade
In the top-left panel of Fig. 6, we report the emitted photons in the $\theta$-$\varepsilon_{e}$ space under three curvature radii at the time $t=300$ ns. Unlike the S-type cascade, the low-energy $e^{\pm}$ pairs created here are accelerated to higher energies by the parallel electric field. Consequently, in addition to CR photons from initial electrons ($\theta<10^{-8}$) and SR photons from created $e^{\pm}$ pairs ($\theta>10^{-4}$), we observe additional SR photons emitted by the accelerated $e^{\pm}$ pairs with intermediate pitch angles ($10^{-8}<\theta<10^{-4}$). As in the S-type cascade, super-SR photons are notably absent in all cases.
@@ 4.3 Photon polarization effect on pair production
Figure 7: Left panels: the temporal evolution of the relative difference of the positron number $\Delta_{+}=(N_{+}-\overline{N}_{+})/\overline{N}_{+}$, where $N_{+}$ and $\overline{N}_{+}$ are the number of generated positrons with and without considering the photon polarization, respectively. Right panels: the QED parameter $\chi_{\gamma}$ of decayed photons with and without the photon polarization. Top panel: S-type cascade with $\varepsilon_{0}=3$ TeV, and the other parameters as in Fig. 5. Bottom panel: A-type cascade with $\rho_{0}=10$ km, and the other parameters as in Fig. 6.
@@
In this subsection, we investigate the influence of photon polarization on pair production. According to Eq. (2), the photon polarization can affect the pair production rate by as much as 30%. When applying Eq. (2), one should pay particular attention to the selected basis for the photon Stokes vector $\bm{\xi}$. The two basis vectors $\bm{e}_{1}$ in Eq. (1) and $\bm{e}_{1}^{\prime}$ of pair production in Eq. (2) are mutually perpendicular for CR. This arises from the fact that the net acceleration direction of $e^{\pm}$ in CR is mainly parallel to the plane of magnetic field lines, while for newly created $e^{\pm}$ pairs, the acceleration direction is mainly perpendicular to it. Consequently, this results in $\xi_{3}\approx-\xi_{3}^{\prime}$. From Eq. (2), the specific polarization of CR photons leads to a higher pair production rate compared to unpolarized photons. This finding starkly contrasts with results obtained in laser fields, where the photon polarization effect was observed to suppress pair production (Wan et al., 2020; Seipt et al., 2021; Song et al., 2021b). In fact, in laser fields, highly polarized photons are emitted via an SR-like mechanism, leading to $\xi_{3}\approx\xi_{3}^{\prime}$.
@@
The relative difference in the positron number $\Delta_{+}=(N_{+}-\overline{N}_{+})/\overline{N}_{+}$ between polarized and unpolarized results is reported in Fig. 7, where $N_{+}$ and $\overline{N}_{+}$ are the number of generated positrons with and without considering the photon polarization, respectively. Figure 7 reveals that the number of created positrons is enhanced by more than 10% at the initial stage of the cascade due to the high polarization of CR photons. In the S-type cascade, at later times the relative difference narrows to less than 5%, even though the polarization of CR photons remains high. The comparison of the $\chi_{\gamma}$ of decayed photons with versus without photon polarization (see the right column of Fig. 7) reveals that for the unpolarized case, $\chi_{\gamma}$ at photon decay is slightly larger than in the polarized case. In fact, unpolarized photons with the same energy as polarized photons posses smaller decay rates and travel longer distances, thus acquiring larger pitch angles, and consequently slightly larger $\chi_{\gamma}$ than polarized photons before decaying into pairs. In the A-type cascade, photon polarization has a similar effect on pair production, as illustrated in the bottom panel of Fig. 7. The relative difference $\Delta_{+}$ can exceed 10% at the initial stage and is also reduced to about 5% with the development of the QED cascade. Similar to the S-type cascade, a slight difference in $\chi_{\gamma}$ is observed between polarized and unpolarized results. At a later stage, a key factor to the reduction in the pair yield difference between polarized and unpolarized photons is the domination of unpolarized SR photons over polarized CR photons.
@@
Figure 8: From top to bottom panel: Spatial distributions of the parallel electric field $E_{\parallel}$ and the normalized positron density $n_{+}/|n_{\rm GJ}|$. Emitted $\gamma$ photons in the $\theta$-$\varepsilon_{e}$ space. Temporal evolution of the photon polarization $\overline{\xi}_{3}$ in relation to the photon energy $\varepsilon_{\gamma}$.
@@ 4.4 Polarized QED-PIC simulations
In our final set of simulations, we employ first-principles 1D (1D geometry but fully 3D particle momentum) QED-PIC simulations in the corotating frame using the YUNIC code (Song et al., 2021a). YUNIC is a relativistic PIC code that includes polarized QED modules and has been previously utilized for studying polarized QED cascades in laser-plasma interactions (Song et al., 2021b; Song et al., 2022). Notably, the electric field in these simulations is inhomogeneous and self-consistently solved. This is a crucial distinction from the single-particle simulations, where the electric field is not evolving. Here the electric field is excited by the outflow of charged $e^{\pm}$ plasmas and subsequently screened by the created dense $e^{\pm}$ plasmas, all of which can be captured by the QED-PIC simulation. Our simulation setup is similar to the pioneering work (Timokhin, 2010). There are several improvements in our QED-PIC code, in which a standard Monte-Carlo QED method developed in recent years by the laser-plasma community is adopted (Elkina et al., 2011; Gonoskov et al., 2015; Ridgers et al., 2014; Tamburini et al., 2017; Montefiori & Tamburini, 2023), and enriched with the polarization effects introduced in section 2. The magnetic field has a strength of $B_{0}=4\times 10^{11}$ G and a curvature radius of $\rho_{0}=10$ km.
@@
With the propagation of the initial flow of $e^{\pm}$ outside the computational box, a parallel electric field is excited, linearly increasing towards the pulsar surface (left boundary) to accelerate $e^{\pm}$ to high energies (see the top panel of Fig. 8). When the QED cascade starts, the density of created $e^{\pm}$ plasma near the pulsar surface can reach 30 times the GJ density at $t=800$ ns, as shown in the second panel from the top of Fig. 8. As expected, the QED cascade creates large amounts of charges that screen the electric field and close the vacuum gap. The $\theta$-$\varepsilon_{e}$ space of photons at different instants is displayed in the third panel from top of Fig. 8. Similar to the single-particle simulation results, CR photons have high energies and small pitch angles, while SR photons exhibit opposite features with lower energies and larger pitch angles than CR photons. The transition from primarily having CR photons to SR photons becoming dominant over time is evident. The average photon polarization $\overline{\xi}_{3}$ as a function of the emitted photon energy is shown in the bottom panel of Fig. 8, with the net polarization of low-energy photons gradually reducing due to the growth of the number of weakly polarized SR photons. An additional insight from the PIC simulation is that high-energy photons are also depolarized at a later stage. This is due to the strong induced electric field, which can accelerate a substantial fraction of secondary $e^{\pm}$ to high energies. Before the electric field is screened, these accelerated $e^{\pm}$ emit high-energy photons while remaining in the SR regime. Despite their high energy, these SR photons are comparatively weakly polarized compared to CR photons. With the inclusion of photon polarization effects, we observe an enhancement of approximately 5% in the $e^{\pm}$ yield compared to the unpolarized results.
@@ Acknowledgements
This work was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (Grants No. XDA25010300 and No. XDA25050300), the National Key R&D Program of China (2021YFA1601700).